# Supernova-induced binary-interaction-powered supernovae: a model for SN2022jli

### v1

@@ 4.1 Super-Eddington accretion
[…]
At such high accretion rates, neutrino emission may start to dominate the cooling in the inner parts of the flow, enabling hypercritical accretion (e.g. Chevalier, 1993, 1996; Fryer et al., 1996). The accretion rate range for neutrino-dominated hypercritical steady accretion onto NSs is thought to be around $10^{-4}~{}M_{\odot}~{}\mathrm{yr}^{-1}\lesssim\dot{M}\lesssim 0.1~{}M_{\odot}~{}\mathrm{yr}^{-1}$, consistent with the accretion rate required to power the luminosity of SN2022jli. In this scenario, the accretion flow is capable of forming a neutrino-cooled disk, which could power bipolar outflows, e.g. via the Blandford-Payne mechanism (Blandford & Payne, 1982), necessary to avoid excessive negative feedback.
@@ 4.2 $\sim$50 d delay to the second peak
In our synthetic light curves, the accretion feedback occurs straight after the SN as the timescale of the companion inflation is on the dynamical timescale. As a result, the accretion feedback kicks in almost immediately, which contradicts with the $\sim 50~{}\mathrm{d}$ delay to the second peak in SN2022jli. Previous studies have shown that the first part of the light curve is consistent with being powered by ${}^{56}\mathrm{Ni}$ decay and can be modelled as a typical type Ic SN. In our hydrodynamic simulations, we have not included the contribution of the SN ejecta, which will greatly contribute to the light at early times but also could be optically thick enough to hide the radiation from the accretion feedback. The $\sim 50~{}\mathrm{d}$ delay to the second peak could therefore correspond to the time it takes for the SN ejecta to become optically thin and expose the interior binary.
@@
[…]
@@
×$\times\left(\frac{M_{\mathrm{ej}}}{M_{\odot}}\right)^{\frac{1}{2}}\left(\frac{v_{\mathrm{ej}}}{10^{4}~{}\mathrm{km~{}s^{-1}}}\right)^{-1},$ (12)
which is consistent with the $\sim 50~{}\mathrm{d}$ delay at the order of magnitude level. Indeed, typical stripped-envelope SNe reach the nebular phase within $\sim 100~{}\mathrm{d}$, roughly consistent with this estimate.
@@ 4.3 Rapid drop at $\sim$250 d
There are several possible explanations for the rapid drop at $\sim 250~{}\mathrm{d}$. First, it could be due to rapid formation of dust, which enshrouds the emitting region. Evidence of dust formation has been reported around the time of the drop (Cartier et al., 2024). Given that the SN ejecta are very C-rich (it is a type Ic SN), it is a perfect site for dust formation. As soon as the ejecta adiabatically cool below a threshold temperature, CO molecules can form, which further enhances the cooling through molecular lines. It then quickly reaches the dust condensation temperature to form dust and enshroud the inner binary. In this case, we may expect that there is continuous energy injection from the embedded binary and the radiation is reprocessed into infrared bands. Although in SN2022jli the infrared flux does increase after the optical drop ($\sim 400~{}\mathrm{d}$), it does not make up for the decrease in optical flux, indicating a drop in the energy injection is also required.
@@
×$\times\left(\frac{M_{\mathrm{NS}}}{1.4~{}M_{\odot}}\right)^{-\frac{2}{3}}\left(\frac{R_{\mathrm{NS}}}{10~{}\mathrm{km}}\right)^{5},$ (13)
@@ 4.4 Gamma-ray detection
In our accretion-powered scenario, the $\gamma$-rays could be generated by a relativistic jet launched from the NS (Figure 13). High-energy ($\sim$100 MeV–TeV) $\gamma$-rays have been observed in the past in stellar mass binaries like LS I$+61^{\circ}303$ (Abdo et al., 2009; Albert et al., 2009). The compact object in this system is most likely a NS (Weng et al., 2022), indicating that NSs are capable of creating GeV $\gamma$-ray emission. While magnetar-like behaviour has been reported (Torres et al., 2012; Weng et al., 2022), bipolar jets have also been detected in radio (Massi et al., 2004, 2012), complicating the identification of the $\gamma$-ray source in LS I$+61^{\circ}303$.
Both mechanisms generate $\gamma$-rays at the shock between the NS outflow (magnetar wind/accretion-powered jet) and the surrounding material, which in this case is the inflated envelope of the companion. The optical depth of the envelope should decline under $\tau\lesssim 1$ for the $\gamma$-rays to emerge. For GeV photons, the main source of opacity is Compton scattering in the Klein-Nishina regime. The Klein-Nishina cross section is approximately $\sigma_{\mathrm{KN}}\approx 3\sigma_{\mathrm{T}}m_{e}c^{2}/(8\epsilon_{\gamma})$ (Lang, 1980; Barkov & Komissarov, 2011), where $\sigma_{\mathrm{T}}$ is the Thomson scattering cross section, $m_{e}$ is the electron mass, $c$ the speed of light and $\epsilon_{\gamma}$ the photon energy. Following Eq. (12), we can estimate the $\gamma$-ray emergence time due to Klein-Nishina scattering as
$t_{\gamma,e}\gtrsim 1.3~{}$ $\mathrm{d}\left(\frac{f}{0.1}\right)^{\frac{1}{2}}\left(\frac{\epsilon_{\gamma}}{1~{}\mathrm{GeV}}\right)^{-\frac{1}{2}}$
×$\times\left(\frac{M_{\mathrm{ej}}}{M_{\odot}}\right)^{\frac{1}{2}}\left(\frac{v_{\mathrm{ej}}}{10^{4}~{}\mathrm{km~{}s^{-1}}}\right)^{-1}.$ (14)
The ejecta would thus become transparent to GeV photons within days.
@@
below which the GeV photons will be absorbed by the keV photons. In our simulations, the polar shock front lies roughly around $R_{\mathrm{sh}}\sim 2000~{}R_{\odot}$ at $\sim 40~{}\mathrm{d}$ (see the outer edge of the polar $v\sim 10^{10}~{}\mathrm{cm~{}s}^{-1}$ component in Figures 10 & 11) and advances linearly in time. If we assume that the X-ray luminosity at the shock is equal to the observed optical luminosity, the absorption radius declines over time. The crossover of the increasing $R_{\mathrm{sh}}$ and decreasing $R_{\mathrm{abs}}$ occurs in a few months, not too far from the $\sim 200~{}\mathrm{d}$ delay in SN2022jli. See Figure 13 for a schematic diagram of the $\gamma$-ray emission in our model.
@@
[…]
@@ 4.7 Future evolution
If the companion star is an OB-type star ($M_{2}\gtrsim 8~{}M_{\odot}$), the system may immediately transition into an eccentric wind-fed high-mass X-ray binary similar to GX 301-2 (Sato et al., 1986). Given that there was a likely mass transfer phase preceding the SN that stripped the hydrogen and helium layers, the companion star could have accreted a substantial amount of mass and angular momentum. Therefore, the companion may eventually evolve into a rapidly-rotating Oe/Be star, turning the system into a Be X-ray binary like LS I$+61^{\circ}303$. This system shares many properties with the SN2022jli system, such as a similar orbital period ($P_{\mathrm{orb}}\sim 26.5~{}\mathrm{d}$; Taylor & Gregory, 1982), eccentricity ($e\sim 0.54$; Aragona et al., 2009) and high-energy $\gamma$-ray emission (Abdo et al., 2009; Albert et al., 2009). In this case, SN2022jli was marking the birth of a high-mass X-ray binary, with an enhanced accretion phase straight after the SN.
@@ 5 Summary and conclusion
Based on our parameter study, a high-eccentricity ($e\gtrsim 0.7$) orbit with geometrically confined feedback is required to reach mass accretion rates compatible with the luminosity of SN2022jli. In addition, the viewing angle should be close to edge-on in order to match the relatively small undulation amplitudes of SN2022jli. Higher viewing angle models are reminiscent of more strongly outbursting SNe like SN2022mop or SN2009ip. On the other hand, the companion mass does not seem to influence the accretion rate significantly.
@@
This work was supported by software support resources awarded under the Astronomy Data and Computing Services (ADACS) Merit Allocation Program. ADACS is funded from the Astronomy National Collaborative Research Infrastructure Strategy (NCRIS) allocation provided by the Australian Government and managed by Astronomy Australia Limited (AAL). This work was partially performed on the OzSTAR national facility at Swinburne University of Technology. The OzSTAR program receives funding in part from the Astronomy NCRIS allocation provided by the Australian Government. Part of the simulations were performed on the supercomputer Gadi with the assistance of resources from the National Computational Infrastructure (NCI Australia), an NCRIS enabled capability supported by the Australian Government. Part of the simulations were also performed on the HOKUSAI computing facility in RIKEN. This work was initiated in part at Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452. S.N. is supported by JSPS Grant-in-Aid Scientific Research (KAKENHI) (A), Grant Number JP25H00675, and by JST ASPIRE Program “RIKEN-Berkeley mathematical quantum science initiative”.
@@ Appendix A Approximate opacity formula
×$\times\left(1+\left(\frac{T}{4.5\times 10^{8}~{}\mathrm{K}}\right)^{0.86}\right)^{-1}~{}\mathrm{cm^{2}~{}g^{-1}},$ (A2)
@@
Figure 15: Opacity as a function of density and temperature computed from Eq. (A1) with $\kappa_{\mathrm{floor}}=0$. Overplotted are the density and temperature of each cell in our simulations from the two snapshots used in Figures 10 & 11.

### v2

@@ 4.1 Super-Eddington accretion
In our simulations, we do not place a limit on the accretion rate onto the NS. Based on the luminosity of the second part of the light curve, the NS needs to create a feedback power of about $\sim 10^{4}~L_{\mathrm{Edd}}$, where $L_{\mathrm{Edd}}$ is the Eddington luminosity. This exceeds the luminosity of the brightest ultra-luminous X-ray pulsar (NGC 5907 ULX-1; Israel et al., 2017) by an order of magnitude. It has been suggested that such high apparent luminosities can be achieved if the radiation is sufficiently geometrically beamed (King & Lasota, 2024). However, in this case the actual feedback power is much less than the isotropic equivalent luminosity of the observed light and thus cannot supply the required energy for our model. Recent general relativistic magneto-hydrodynamics simulations of super-Eddington accretion onto NSs demonstrate that it is possible at relatively low magnetic field strengths $B\lesssim 10^{11}$ G (Inoue et al., 2023; Kayanikhoo et al., 2025) to successfully produce radiative and kinetic outflow luminosities reaching $\sim 500$$1000~L_{\mathrm{Edd}}$, close to what is required for NGC 5907 ULX-1. Such high accretion rates are sustained due to the break in spherical symmetry as the accretion flow is channeled along magnetic field lines towards the magnetic poles. Radiative losses from the sides of the accretion columns cool the matter enough to allow accretion at highly super-Eddington rates.
At such high accretion rates, neutrino emission may start to dominate the cooling in the inner parts of the flow, enabling hypercritical accretion (e.g. Chevalier, 1993, 1996; Fryer et al., 1996). The accretion rate range for neutrino-dominated hypercritical steady accretion onto NSs is thought to be around $10^{-4}~M_{\odot}~\mathrm{yr}^{-1}\lesssim\dot{M}\lesssim 0.1~M_{\odot}~\mathrm{yr}^{-1}$, consistent with the accretion rate required to power the luminosity of SN2022jli. In this scenario, the accretion flow is capable of forming a neutrino-cooled disk, which could power bipolar outflows, e.g. via the Blandford-Payne mechanism (Blandford & Payne, 1982), necessary to avoid excessive negative feedback. However, in this regime it is also expected that a hydrostatic atmosphere is formed (Chevalier, 1993; Combi et al., 2025), significantly lowering the accretion feedback luminosity. We note that these neutrino-cooled models tend to ignore the NS magnetic field and thus the full picture of neutrino-cooled, magnetically channeled accretion is not understood.
@@ 4.2 $\sim$50 d delay to the second peak
In our synthetic light curves, the accretion feedback occurs straight after the SN as the timescale of the companion inflation is on the dynamical timescale. As a result, the accretion feedback kicks in almost immediately, creating noticeable features in the early light curve. Previous studies have shown that the first part of the light curve is consistent with being powered by $^{56}\mathrm{Ni}$ decay and can be modelled as a typical type Ic SN. In our hydrodynamic simulations, we have not included the contribution of the SN ejecta, which will greatly contribute to the light at early times but also could be optically thick enough to hide the radiation from the accretion feedback. The $\sim 50~\mathrm{d}$ delay to the second peak could therefore correspond to the time it takes for the SN ejecta to become optically thin and expose the interior binary.
@@
$\tau_{\mathrm{SN}}=\int_{0}^{\infty}\kappa\rho_{\mathrm{ej}}dr=f\bar{\kappa}\frac{M_{\mathrm{ej}}}{R_{\mathrm{ej}}^{3}}R_{\mathrm{ej}},$ (14)
@@
$\times\left(\frac{M_{\mathrm{ej}}}{M_{\odot}}\right)^{\frac{1}{2}}\left(\frac{v_{\mathrm{ej}}}{10^{4}~\mathrm{km~s^{-1}}}\right)^{-1},$ (15)
which is consistent with the $\sim 50~\mathrm{d}$ delay at the order of magnitude level. Indeed, typical stripped-envelope SNe reach the nebular phase within $\sim 100~\mathrm{d}$, roughly consistent with this estimate. For the case of SN2022jli, nebular features start to emerge in the spectrum at around the second peak of the light curve (Cartier et al., 2024). This may indicate that the SN ejecta is sufficiently optically thin at this point but the spectrum does not become completely nebular due to the emission from the underlying binary.
[…]
@@ 4.3 Rapid drop at $\sim$250 d
There are several possible explanations for the rapid drop at $\sim 250~\mathrm{d}$. First, it could be due to rapid formation of dust, which enshrouds the emitting region. Early dust formation has been detected in some known type Ic SNe (Rho et al., 2021; Ravi et al., 2023), and evidence of dust formation has been reported for SN2022jli around the time of the drop (Cartier et al., 2024). Given that the SN ejecta are very C-rich (it is a type Ic SN), it is a good site for dust formation. As soon as the ejecta adiabatically cool below a threshold temperature, CO molecules can form, which further enhances the cooling through molecular lines. It then quickly reaches the dust condensation temperature to form dust and enshroud the inner binary. If dust was the reason for the drop, we may expect that there is continuous energy injection from the embedded binary and the radiation is reprocessed into infrared bands. Although in SN2022jli the infrared flux does increase after the optical drop ($\sim 400~\mathrm{d}$), it does not make up for the decrease in optical flux, indicating a drop in the energy injection is also required.
@@
$\times\left(\frac{M_{\mathrm{NS}}}{1.4~M_{\odot}}\right)^{-\frac{2}{3}}\left(\frac{R_{\mathrm{NS}}}{10~\mathrm{km}}\right)^{5},$ (16)
@@ 4.4 Gamma-ray detection
[…]
Both mechanisms (magnetar or relativistic jets) generate $\gamma$-rays at the shock between the NS outflow (magnetar wind/accretion-powered jet) and the surrounding material, which in this case is the inflated envelope of the companion. The optical depth of the envelope should decline under $\tau\lesssim 1$ for the $\gamma$-rays to emerge. The $\gamma$-rays were only detected at $\sim 200~\mathrm{d}$, meaning that it must have been obscured before that. For GeV photons, the main source of opacity is the Bethe-Heitler pair production process (Vurm & Metzger, 2021; Lu et al., 2025). The effective opacity is given by (Zdziarski & Svensson, 1989; Vurm & Metzger, 2021)
$\kappa_{\gamma p}=\frac{21}{2\pi}\alpha_{\mathrm{fs}}\left(\ln{\left[\frac{2\epsilon_{\gamma}}{m_{e}c^{2}}\right]}-\frac{109}{42}\right)\kappa_{\mathrm{T}},$ (17)
where $\alpha_{\mathrm{fs}}\sim 1/137$ is the fine structure constant, $\epsilon_{\gamma}$ is the photon energy, $m_{e}$ is the electron mass, $c$ the speed of light and $\kappa_{\mathrm{T}}$ is the Thomson scattering opacity. For GeV photons, this becomes $\kappa_{\gamma p}\sim 0.03~\mathrm{g~cm^{-2}}$. Following Eq. (15), we can estimate the $\gamma$-ray emergence time as
$t_{\gamma,\mathrm{ej}}\gtrsim 30$~ $\mathrm{d}\left(\frac{f}{0.1}\right)^{\frac{1}{2}}\left(\frac{M_{\mathrm{ej}}}{M_{\odot}}\right)^{\frac{1}{2}}\left(\frac{v_{\mathrm{ej}}}{10^{4}~\mathrm{km~s^{-1}}}\right)^{-1}.$ (18)
The ejecta would thus become transparent to GeV photons within months, although it may be compatible with $\sim 200~\mathrm{d}$ if the ejecta mass was larger and the associated ejecta velocity was slower.
@@
below which the GeV photons will be absorbed by the keV photons. In our simulations, the polar shock front lies roughly around $R_{\mathrm{sh}}\sim 2000~R_{\odot}$ at $\sim 50~\mathrm{d}$ (see the outer edge of the polar $v\sim 10^{10}~\mathrm{cm~s}^{-1}$ component in Figures 13 & 14) and advances linearly in time. If we assume that the X-ray luminosity at the shock is equal to the observed optical luminosity, the absorption radius declines over time. The crossover of the increasing $R_{\mathrm{sh}}$ and decreasing $R_{\mathrm{abs}}$ occurs in a few months, not too far from the $\sim 200~\mathrm{d}$ delay in SN2022jli. See Figure 16 for a schematic diagram of the $\gamma$-ray emission in our model.
@@
It is still questionable whether the cyclic pattern they identify are due to the orbital motion of the companion. The velocity profile is rather complex with multiple peaks and the wings spread out to $|v_{\mathrm{H}\alpha}|\sim 4000~\mathrm{km~s}^{-1}$. In fact, there are no obvious peaks in the orbital velocity range inferred from the flux-weighted centroid ($|v_{\mathrm{H}\alpha}|<600~\mathrm{km~s}^{-1}$; Chen et al., 2024). Instead, there appears to be a steady broad component with a maximum at around $v_{\mathrm{H}\alpha}\sim+1000$$2000~\mathrm{km~s}^{-1}$ and another component at $v_{\mathrm{H}\alpha}\sim-1000~\mathrm{km~s}^{-1}$ that strengthens around the undulation peaks (Cartier et al., 2024).
@@ 4.7 Future evolution
If the companion is a lower mass star ($M_{2}\sim 2$$5~M_{\odot}$), the system will soon become dormant, as there will be neither Roche lobe overflow nor sufficiently strong winds to maintain accretion. In this detached phase, it could be observed as an astrometric binary with a dark companion (e.g. Gaia NS1; El-Badry et al., 2024a, b). It will only re-activate once the companion evolves to emit strong winds and become a symbiotic X-ray binary or overflow its Roche lobe and transition into a low/intermediate-mass X-ray binary phase. A system like Circinus X-1 could be a possible end product, which again shares many similarities with SN2022jli such as the orbital period and eccentricity ($P_{\mathrm{orb}}\sim 16.5~\mathrm{d},e\sim 0.45$; Jonker et al., 2007). The current X-ray light curve bears a striking similarity to the SN2022jli optical light curve, with saw-tooth-like modulations (Tominaga et al., 2023). It is also known to launch relativistic jets (Fender et al., 2004). It is possible that SN2022jli was a Cir X-1-like system embedded in an inflated envelope, which eventually transitioned out of the accretion phase but will return in the future.
@@ 5 Summary and conclusion
Based on our parameter study, a high-eccentricity ($e\gtrsim 0.8$) orbit with geometrically confined feedback is required to reach mass accretion rates compatible with the luminosity of SN2022jli. The viewing angle dependence is insignificant at these high eccentricities, but become increasingly important at $e\lesssim 0.7$. Higher viewing angle models with lower eccentricity are reminiscent of more strongly outbursting SNe like SN2022mop or SN2009ip. On the other hand, the companion mass does not seem to influence the accretion rate significantly.
@@
The authors thank the anonymous referee for constructive comments that significantly improved the quality of the work. RH thanks Akihiro Inoue for useful discussions on super-Eddington accretion. This work was supported by software support resources awarded under the Astronomy Data and Computing Services (ADACS) Merit Allocation Program. ADACS is funded from the Astronomy National Collaborative Research Infrastructure Strategy (NCRIS) allocation provided by the Australian Government and managed by Astronomy Australia Limited (AAL). This work was partially performed on the OzSTAR national facility at Swinburne University of Technology. The OzSTAR program receives funding in part from the Astronomy NCRIS allocation provided by the Australian Government. Part of the simulations were performed on the supercomputer Gadi with the assistance of resources from the National Computational Infrastructure (NCI Australia), an NCRIS enabled capability supported by the Australian Government. Part of the simulations were also performed on the HOKUSAI computing facility in RIKEN. This work was initiated in part at Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452. S.N. is supported by JSPS Grant-in-Aid Scientific Research (KAKENHI) (A), Grant Number JP25H00675 and (B), Grant Number JP23K25874, and by JST ASPIRE Program “RIKEN-Berkeley mathematical quantum science initiative”.
@@ Appendix A Approximate opacity formula
$\times\left(1+\left(\frac{T}{4.5\times 10^{8}~\mathrm{K}}\right)^{0.86}\right)^{-1}~\mathrm{cm^{2}~g^{-1}},$ (A2)
@@
Figure 18: Opacity as a function of density and temperature computed from Eq. (A1) with $\kappa_{\mathrm{floor}}=0$. Overplotted are the density and temperature of each cell in our simulations from the two snapshots used in Figures 13 & 14.
Appendix B 1D explosion simulation
To roughly represent the density profile of type Ic SN ejecta, we employ a thermal bomb technique to explode a stellar model of a naked CO core. We first evolve a $15~M_{\odot}$ star in MESA, mostly adopting the inlists in the 20M_pre_ms_to_core_collapse test suite. After core He depletion, we reduce the mass down to $3~M_{\odot}$, stripping off the H and He-rich layers. This exposes the CO-rich layers to the surface, representing a type Ic SN progenitor. The star is further evolved up to core collapse. We then map the MESA stellar profile into a 1D spherical coordinate grid in HORMONE, excising the inner $1.5~M_{\odot}$ from the computational domain to represent the proto neutron star. The radial direction is split into 1500 grid points where the grid spacing is increased outwards in a geometrical series and the inner most cell size is $\Delta r_{1}=10^{7}~\mathrm{cm}$. At $t=0$, we inject $E_{\mathrm{inj}}=1.18\times 10^{51}~\mathrm{erg}$ in the innermost 10 cells to drive an explosion. Subtracting off the envelope binding energy, this leads to an ejecta kinetic energy of $E_{\mathrm{exp}}=10^{51}~\mathrm{erg}$. A low-density atmosphere is attached outside the star that has negligible mass compared to the star. A reflective boundary is applied at the inner boundary and an outgoing boundary is applied at the outer boundary.
As soon as the simulation is started, a strong outgoing shock is formed around the energy-added cells. We follow the evolution until the shock breaks out of the stellar surface and reaches a state of homologous expansion. For the red curves in Figures 13 & 14, we take the final snapshot and appropriately rescale the density profile assuming homologous expansion.